The product of two numbers is 4 square root of 3. Find the numbers so that the
sum 𝑆 S of the square of one and the cube of the other is as small as
possible.




1
Expert's answer
2021-06-14T18:09:30-0400

Given ab = 4*3(1/2) , a = 4*3(1/2)/b

Given f(a,b) = a2 + b3 = (48/b2)+b3

Differentiating both sides with respect to b

f'(a,b) = 3b2 - (96/b3)

For critical points, f'(a,b) = 0

3b2 = 96/b3

b5 = 32

b = 2

Again Differentiating both sides with respect to b

f"(a,b) = 6b + 384/b4 > 0

At b = 2, f(a,b) has local minimum.

So, if b = 2, a = 2*3(1/2)